# Adaptive Flip Graph Algorithm for Matrix Multiplication

### 28 Dec 2023 v1

@@ 4 Results
4.1 Effectiveness of Plus Transition Escaping Non-Reduction States
@@
Figure 1: Dependence of the minimum rank found in the search on the number of iterations in the $(4,4,4)$-matrix multiplication scheme. Results for the flip graph algorithm with the plus transition are marked with circles and those without the plus transition are marked with triangles. Error bars represent the standard error obtained by 10 independent runs.
4.2 Acceleration by Edge Constraints in Large Matrix Multiplications
@@
Figure 2: Dependence of the minimum rank found in the search on the number of iterations in the $(5,5,5)$-matrix multiplication scheme. Results for the flip graph algorithm with the edge constraints are marked with circles and those without the edge constraints are marked with triangles. Error bars represent the standard error obtained by 10 independent runs.
4.3 Results of Adaptive Flip Graph Algorithm for various matrix multiplications
We have confirmed that both plus transitions and edge constraints improve the efficiency of the flip graph algorithm from different perspectives. The adaptive flip graph algorithm, incorporating these two ideas, was applied to the $(n,m,p)$-matrix multiplications in characteristic 2. For the results of the adaptive flip graph algorithm, the plus flag was set to $L=5\times 10^{3}$ for all edge constraints. In the case of the maximal case of the $(5,5,5)$-matrix multiplication, the total number of iterations of edge constraints was set as $\mathcal{T}=\{5\times 10^{7}$, $5\times 10^{7}$, $5\times 10^{7}$, $5\times 10^{7}$, $10^{8}$, $1.5\times 10^{8}$, $2.5\times 10^{8}$, $7.5\times 10^{8}$, $1.5\times 10^{9}$, $3\times 10^{9}\}$ and the edge constraints were repeated $10^{2}$ times, starting from the previous edge constraints scheme with lowest scheme rank. As shown in Table 1, the adaptive flip graph algorithm achieved updated minimum ranks for the $(4,5,5)$ and $(5,5,5)$-matrix multiplication schemes, while reproducing the previously known minimum ranks for other smaller matrix multiplication schemes. The schemes we found for $(4,5,5)$ and $(5,5,5)$-matrix multiplications are explicitly shown in Appendix A and B, respectively.
Table 1: Comparison of the best-known rank and the rank found by the adaptive flip graph algorithm in characteristic 2 for various matrix multiplications, denoted as “Adap. Flip”. The table includes results obtained by the Alpha Tensor based reinforcement learning (Fawzi et al., 2022) denoted as “AT”, those obtained by the flip graph algorithm starting from the standard classical matrix multiplication algorithms (Kauers and Moosbauer, 2023) denoted as “Flip” and those obtained by the flip graph algorithm starting from the “AT” results examined only in the $(5,5,5)$-matrix multiplication denoted as “FwAT” (Kauers and Moosbauer, 2023).
@@ 4.3 Results of Adaptive Flip Graph Algorithm for various matrix multiplications
$(2,2,2)$ Strassen (1969) 7 7 7 7
$(2,2,3)$ Hopcroft and Kerr (1971) 11 11 11 11
$(2,2,4)$ Hopcroft and Kerr (1971) 14 14 14 14
$(2,3,3)$ Hopcroft and Kerr (1971) 15 15 15 15
$(2,2,5)$ Hopcroft and Kerr (1971) 18 18 18 18
$(2,3,4)$ Hopcroft and Kerr (1971) 20 20 20 20
$(3,3,3)$ Laderman (1976) 23 23 23 23
$(2,3,5)$ Hopcroft and Kerr (1971) 25 25 25 25
$(2,4,4)$ Hopcroft and Kerr (1971) 26 26 26 26
$(3,3,4)$ Smirnov (2013) 29 29 29 29
$(2,4,5)$ Hopcroft and Kerr (1971) 33 33 33 33
$(3,3,5)$ Smirnov (2013) 36 36 36 36
$(3,4,4)$ Smirnov (2013) 38 38 38 38
$(2,5,5)$ Hopcroft and Kerr (1971) 40 40 40 40
$(3,4,5)$ Fawzi et al. (2022) 47 47 47 47
$(4,4,4)$ Fawzi et al. (2022) 47 47 47 47
$(3,5,5)$ Sedoglavic and Smirnov (2021) 58 58 58 58
$(4,4,5)$ Kauers and Moosbauer (2023) 60 63 60 60
$(4,5,5)$ Fawzi et al. (2022) 76 76 76 $\bm{73}$
$(5,5,5)$ Kauers and Moosbauer (2023) 95 96 97 95 $\bm{94}$
5 Discussion and Summary
@@
We compare our study with previous studies from the perspective of computational resources. All numerical experiments in this study were performed on an Intel(R) Core(TM) i7-8650U CPU 1.90GHz 2.11 GHz, single CPU thread. While the previous reinforcement learning study using the Alpha Tensor (Fawzi et al., 2022) employed a 64-core TPU and the original flip graph algorithm (Kauers and Moosbauer, 2023) required massive parallel computational resources, our method demonstrated high efficiency with only a single CPU thread.
Moreover, we contrast our approach with the method presented in the original flip graph algorithm (Kauers and Moosbauer, 2023) with respect to the initial schemes. The previous study found the scheme with rank 95 for the $(5,5,5)$-matrix multiplication, using the Alpha Tensor results as the initial scheme. In contrast, when using the standard classical matrix multiplication algorithm as the initial scheme, the best scheme found was the scheme of rank 97. It should be emphasized that our method successfully found the scheme of rank 94 using the standard classical matrix multiplication algorithm as the initial scheme.

### 16 Mar 2024 v2

@@ 6. Results
6.1. Effectiveness of Plus Transition Escaping Non-Reduction States
@@
Figure 1. Dependence of the minimum rank found in the search on the number of iterations in the $(4,4,4)$-matrix multiplication scheme. Results for the flip graph algorithm with the plus transition are marked with circles and those without the plus transition are marked with triangles. Error bars represent the standard error obtained by 10 independent runs.
6.2. Acceleration by Edge Constraints in Large Matrix Multiplications
@@
Figure 2. Dependence of the minimum rank found in the search on the number of iterations in the $(5,5,5)$-matrix multiplication scheme. Results for the flip graph algorithm with the edge constraints are marked with circles and those without the edge constraints are marked with triangles. Error bars represent the standard error obtained by 10 independent runs.
6.3. Results of Adaptive Flip Graph Algorithm for Various Matrix Multiplications
We have confirmed that both plus transitions and edge constraints improve the efficiency of the flip graph algorithm from different perspectives. The adaptive flip graph algorithm, incorporating these two ideas, was applied to the $(n,m,p)$-matrix multiplications in characteristic 2. For the results of the adaptive flip graph algorithm, the plus flag was set to $L=5\times 10^{3}$ for all edge constraints. In the case of the maximal case of the $(5,5,5)$-matrix multiplication, the total number of iterations of edge constraints was set as $\mathcal{T}=\{5\times 10^{7}$, $5\times 10^{7}$, $5\times 10^{7}$, $5\times 10^{7}$, $10^{8}$, $1.5\times 10^{8}$, $2.5\times 10^{8}$, $7.5\times 10^{8}$, $1.5\times 10^{9}$, $3\times 10^{9}\}$ and the edge constraints were repeated $10^{2}$ times, starting from the previous edge constraints scheme with lowest scheme rank. As shown in Table 1, the adaptive flip graph algorithm achieved updated minimum ranks for the $(4,5,5)$ and $(5,5,5)$-matrix multiplication schemes, while reproducing the previously known minimum ranks for other smaller matrix multiplication schemes. The schemes we found for $(4,5,5)$ and $(5,5,5)$-matrix multiplications are explicitly shown in https://github.com/Yamato-Arai/adap.
Table 1. Comparison of the best-known rank and the rank found by the adaptive flip graph algorithm in characteristic 2 for various matrix multiplications, denoted as “Adap. Flip”. The table includes results obtained by the Alpha Tensor-based reinforcement learning (Fawzi2022, 2) denoted as “AT”, those obtained by the flip graph algorithm starting from the standard classical matrix multiplication algorithms (kauers2022flip, 4) denoted as “Flip” and those obtained by the flip graph algorithm starting from the “AT” results examined only in the $(5,5,5)$-matrix multiplication denoted as “FwAT” (kauers2022flip, 4).
@@ 6.3. Results of Adaptive Flip Graph Algorithm for Various Matrix Multiplications
$(2,2,2)$ (strassen1969gaussian, 9) 7 7 7 7
$(2,2,3)$ (hopcroft1971minimizing, 3) 11 11 11 11
$(2,2,4)$ (hopcroft1971minimizing, 3) 14 14 14 14
$(2,3,3)$ (hopcroft1971minimizing, 3) 15 15 15 15
$(2,2,5)$ (hopcroft1971minimizing, 3) 18 18 18 18
$(2,3,4)$ (hopcroft1971minimizing, 3) 20 20 20 20
$(3,3,3)$ (laderman1976noncommutative, 6) 23 23 23 23
$(2,3,5)$ (hopcroft1971minimizing, 3) 25 25 25 25
$(2,4,4)$ (hopcroft1971minimizing, 3) 26 26 26 26
$(3,3,4)$ (smirnov2013bilinear, 8) 29 29 29 29
$(2,4,5)$ (hopcroft1971minimizing, 3) 33 33 33 33
$(3,3,5)$ (smirnov2013bilinear, 8) 36 36 36 36
$(3,4,4)$ (smirnov2013bilinear, 8) 38 38 38 38
$(2,5,5)$ (hopcroft1971minimizing, 3) 40 40 40 40
$(3,4,5)$ (Fawzi2022, 2) 47 47 47 47
$(4,4,4)$ (Fawzi2022, 2) 47 47 47 47
$(3,5,5)$ (sedoglavic2021tensor, 7) 58 58 58 58
$(4,4,5)$ (kauers2022flip, 4) 60 63 60 60
$(4,5,5)$ (Fawzi2022, 2) 76 76 76 $\bm{73}$
$(5,5,5)$ (kauers2022flip, 4) 95 96 97 95 $\bm{94}$
7. Discussion and Summary
@@
We compare our study with previous studies from the perspective of computational resources. All numerical experiments in this study were performed on an Intel(R) Core(TM) i7-8650U CPU 1.90GHz 2.11 GHz, single CPU thread. While the previous reinforcement learning study using the Alpha Tensor (Fawzi2022, 2) employed a 64-core TPU and the original flip graph algorithm (kauers2022flip, 4) required massive parallel computational resources, our method demonstrated high efficiency with only a single CPU thread.
Moreover, we contrast our approach with the method presented in the original flip graph algorithm (kauers2022flip, 4) with respect to the initial schemes. The previous study found the scheme with rank 95 for the $(5,5,5)$-matrix multiplication, using the Alpha Tensor results as the initial scheme. In contrast, when using the standard classical matrix multiplication algorithm as the initial scheme, the best scheme found was the scheme of rank 97. It should be emphasized that our method successfully found the scheme of rank 94 using the standard classical matrix multiplication algorithm as the initial scheme.